Circular No. 8619

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 T4 (SWAN) R. D. Matson, Newport Coast, California; and M. Mattiazzo, Walla-

roo, South Australia, have independently reported the presence of a comet in SOHO SWAN images during the first half of October (the object appearing rather faint in the ultraviolet images). The SWAN positions below are from Matson. E. J. Christensen reports that Catalina Sky Survey (0.68-m Schmidt telescope) images on Oct. 22.1 UT show a moderately condensed, roughly circular coma of diameter ∼ 2′ in three coadded 20second exposures. R. H. McNaught, observing with the 1.0-m f/8 reflector at Siding Spring Observatory on Oct. 22.4, reports a diffuse 1′ coma with a 3′′ (FWHM) central condensation (used for the magnitude estimates below). J. E. McGaha (Tucson, Arizona, 0.62-m f/5.1 reflector) reports that his images on Oct. 23.1 show a fan-shaped coma of size 20′′ × 30′′ toward p.a. 120^{o}. A. Hale (Cloudcroft, New Mexico, 0.41-m reflector) writes that a visual observation on Oct. 23.08 showed a diffuse coma of diameter 1′.5 and total mag 12.1.

2005 UT  α_{2000}  δ_{2000}  Mag. Observer
Oct. 6  15^{h}00.^{m}7  − 0^{o}20′  SWAN
9  15 29.2  − 3 28  _{”}
11  15 47.9  − 5 10  _{”}
13  16 06.9  − 6 43  _{”}
22.06892 16 27 47.^{s}50 − 9 13 25′′.2 12.4 Christensen
22.07064 16 27 47.68 − 9 13 27.0  _{”}
22.07230 16 27 47.91 − 9 13 28.7  _{”}
22.39354 16 28 24.32 − 9 17 53.4 18.1 McNaught
22.39504 16 28 24.50 − 9 17 55.1 18.0  _{”}
22.39654 16 28 24.67 − 9 17 56.3 18.0  _{”}
22.39809 16 28 24.83 − 9 17 57.4 17.9  _{”}
22.39960 16 28 24.99 − 9 17 58.8 17.9  _{”}
22.40110 16 28 25.18 − 9 18 00.0 17.7  _{”}

Additional precise positions, the following preliminary parabolic orbital elements by B. G. Marsden (from 12 precise positions, Oct. 22–23), and an ephemeris appear on MPEC 2005-U19:

T = 2005 Oct. 9.602 TT  ω = 40.^{o}696
Ω = 25.945 }2000.0
q = 0.64699 AU  i = 160.126
2005 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green